### v2

@@ Abstract
Domain shift occurs when training (source) and test (target) data diverge in their distribution. Source-Free Domain Adaptation (SFDA) addresses this domain shift problem, aiming to adopt a trained model on the source domain to the target domain in a scenario where only a well-trained source model and unlabeled target data are available. In this scenario, handling false labels in the target domain is crucial because they negatively impact the model performance. To deal with this problem, we propose to update cluster prototypes (i.e., centroid of each sample cluster) and their structure in the target domain formulated by the source model in online manners. In the feature space, samples in different regions have different pseudo-label distribution characteristics affected by the cluster prototypes, and we adopt distinct training strategies for these samples by defining clean and noisy regions: we selectively train the target with clean pseudo-labels in the clean region, whereas we introduce mix-up inputs representing intermediate features between clean and noisy regions to increase the compactness of the cluster. We conducted extensive experiments on multiple datasets in online/offline SFDA settings, whose results demonstrate that our method, CNG-SFDA, achieves state-of-the-art for most cases.
@@
Source-Free Domain Adaptation Clean-and-Noisy Region Guiding Online Prototype
@@ 1 Introduction
In this work, we introduce Clean-and-Noisy region Guided SFDA, namely CNG-SFDA, a novel SFDA approach that focuses on addressing pseudo-labels of samples near cluster boundaries given the domain knowledge from a source model. We first partition Clean and Noisy regions within clusters formed in the target domain by leveraging online-updating cluster prototype (i.e., the centroid of each cluster) to measure the distance between the prototypes and the unlabeled target samples. Then, as shown in the bottom of Fig. 1, distinct training strategies are adopted for the clean and noisy regions due to their divergent local characteristics.
For the clean region, we conduct selective training using reliable pseudo-labels derived from predictions of the nearest neighborhoods. In the noisy region, we leverage mix-up inputs that represent intermediate features in gray zone bridging clean and noisy regions in Fig. 1, unlike other methods that directly employ or discard the unreliable pseudo-labels for training. To draw these mix-up inputs toward the clean region effectively, we assign a “clean probability” to each sample indicating the likelihood of having clean pseudo-labels and introduce a mixed-clean probability defined from the clean probability of all target samples with clean and noisy, thereby enabling effective training on samples in noisy regions. Finally, in order to generate reliable pseudo-labels, we introduce the contrastive learning framework, aligning the nearest features effectively. Our method utilizes instance-discriminate and class-semantic features by introducing the memory bank [18] and the cluster prototype embedding through contrastive learning.
@@
  • We extensively validate CNG-SFDA on various benchmarks for online/offline SFDA in single-source and multi-source settings, which empirically demonstrate the superiority of our approach to SOTA techniques by a large margin in the challenging scenarios.
@@ 2.2 Online-Offline Source-Free Domain Adaptation
Recently, several methods [31, 51, 4, 25, 64, 33] have been proposed for handling the domain shift in online-offline source-free domain adaptation (SFDA) settings. SHOT [31], as a self-training method, introduces a class centroid used for pseudo-labeling for offline SFDA. Tent [51] optimizes entropy minimization for online SFDA. CoWA-JMDS [28] utilizes the joint model-data structure as sample-wise weights to represent target domain knowledge effectively. CRCo [64] proposes a probability-based similarity between target samples by embedding the source domain class relationship.
@@
Figure 2: Framework of CNG-SFDA Section (a): partition the cluster into ‘Clean’ (close to the cluster prototype) and ‘Noisy’ (far from the cluster prototype) region based on the clean probability distribution. Section (b): training samples in clean regions with pseudo-labels generated from prediction of closest features. Section (c): reducing a gap between clean and noisy region with mix-up weight. Section (d): prototype-aware and instance-aware contrastive learning.
@@ 3 The Proposed Method
In this work, we address Source-Free Domain Adaptation (SFDA) for the image classification task. In SFDA, we are given a source model trained on labeled source data $\mathcal{D}_{s}=\{x_{s}^{i},y_{s}^{i}\}_{i=1}^{N_{s}}$ and unlabeled target data $\mathcal{D}_{t}=\{x_{t}^{i}\}_{i=1}^{N_{t}}$, where $x_{s}^{i}$, $x_{t}^{i}$ represent the input images and $N_{s}$, $N_{t}$ denote the numbers of data in the source and target domains, respectively. We focus on the closed-set source-free domain adaptation where the target domain shares the same $C$ classes as the source domain.
@@ 3.1 Distinguishing Clean and Noisy Regions in Clusters
$\small{\mu}_{k}=\left\|\frac{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)g(x_{t}^{i})}{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)}\right\|_{\ell_{2}},$ (4)
$\small{\sigma}_{k}=\frac{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)(g(x_{t}^{i})-\mu_{k})^{T}(g(x_{t}^{i})-\mu_{k})}{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)}.$ (5)
With the above parameters ${\mu}_{k},{\sigma}_{k}$ estimated from $f(\cdot)$, the Eq. (3) is defined by the probability density function:
@@ 3.2 Training Strategy for Clean and Noisy Regions
In our preliminary study, we observed that the clean and noisy regions within a cluster have divergent local characteristics. In the clean region, the samples, as well as their closest features, have reliable pseudo-labels $\hat{y}_{t}$ generated by Eq. 1, as they are also in the clean regions. On the other hand, the noisy regions contain samples with incorrect $\hat{y}_{t}$, as they are located near or on the cluster boundaries. To deal with this issue, we employ distinct training strategies for $\{x_{t}^{cr}\}_{cr=1}^{N_{cr}}$ and $\{x_{t}^{nr}\}_{nr=1}^{N_{nr}}$.
@@
Additionally, we add a regularization that promotes diverse predictions similar to [4] as:
@@
where $\bar{p}_{t}=\mathbb{E}_{x_{t}\in\mathcal{X}_{t}}\sigma(f(T_{s}(x_{t})))$ and the $L_{div}$ prevents the model from trusting false pseudo-labels.
@@ 3.3 Instance and Prototype-aware Contrastive Learning
Moreover, to learn instance-discriminate features, we leverage queries $q_{t}$ and keys $k_{t}=\bar{g}(T_{\bar{s}}(x_{t}))$ as the positive pair as in [18]. To build the negative pairs $\mathcal{N}_{p}$ of $q_{t}$, we store $k_{t}$ to the memory queue $Q_{k}=\{k_{t}^{i}\}_{i=1}^{M}$ with the length $M$, but $Q_{k}$ may have features which have the same class as the corresponding class $\hat{y}_{t}$ of $q_{t}$. In our work, to eliminate negative pairs with the same class, we utilize the nearest feature information in $Q_{k}$. We exclude the negative pairs when classes $\{\hat{y}_{t}^{n}\}_{n=1}^{N}$ of their $N$ closest features are at least one the same as $\hat{y}_{t}$, as shown in Fig. 2 (d). Our instance-aware contrastive loss reflecting the behavior above is given by:
@@ 4.1 Experimental Setup
4.2 Experimental Results
Table. 2, 4, and 4 show the performance of CNG-SFDA and the compared methods across 3 different datasets in the single-source setting. Table. 6 shows results in the multi-source setting. In each table, $\dagger$ indicates the results that we reproduced from the released source code, and ‘SF’ and ‘On’ denote ‘source-free’ and ‘online’, respectively. All reported results are average accuracies from three seeds.
[…]
@@ 4.2 Experimental Results
Multi-Source Domain Adaptation. Table. 6 presents the performance of Multi-Source Unsupervised Domain Adaptation (MSUDA) on PACS. In MSUDA, we treat each domain as the target domain and train the source model by aggregating all other domains without using domain labels. In a challenge source-free MSUDA, CNG-SFDA achieves the highest accuracy on PACS. SHOT++ [32] produces pseudo-labels of the target based on the centroid of the nearest cluster. In contrast, CNG-SFDA trains pseudo-labels for the target data with different strategies depending on whether the data is in clean or noisy regions. These results demonstrate that CNG-SFDA outperforms prior source-free MSUDA methods in terms of learning pseudo-label.
Method SF On A C P S Avg
SIB [20] 88.9 89.0 98.3 82.2 89.6
T-SVDNET [30] 90.4 90.6 98.5 85.4 91.2
iMSDA[27] 93.7 92.4 98.4 89.2 93.4
Source Only - - 77.9 72.8 95.7 63.5 77.5
$\text{SHOT}^{\dagger}$ [31] 90.7 88.1 98.5 75.4 88.2
$+$ [32] 92.3 89.7 98.8 75.5 89.1
CNG-SFDA 93.7 93.3 98.7 87.8 93.4
Table 5: Classification Accuracy (%) on PACS for Multi-Source Domain Adaptation (ResNet-18 Backbone).
PL $L_{cr}$ $L_{ccp}$ $L_{div}$ $L_{inst}$ $L_{prt}$ VisDA-C DomainNet PACS
84.9 58.5 75.7
88.0 65.8 78.8
89.6 67.4 79.7
89.0 67.7 80.6
90.1 69.7 81.1
90.5 70.4 82.4
Table 6: The effectiveness of each component across all datasets is validated by classification accuracy (%).
@@ 4.3 Analysis and Discussion
Figure 3: Feature distribution of clean and noisy samples in target domain. Here, ‘clean’ and ‘noisy’ refer to cases where the Ground-Truth (GT) and the Pseudo-Label (PL) of the samples match and mismatch, respectively. Circle and triangular denote clean and noisy samples predicted by CNG-SFDA, respectively. For clean samples (circle), each color represent the ground-truth class. For noisy samples (triangular), pink and black triangular represent cases where CNG-SFDA correctly (i.e., Prediction: Noisy, Ground-Truth: Noisy) and incorrectly (i.e., Prediction: Noisy, Ground-Truth: Clean) identifies noisy data, respectively. (a), (c): feature distribution before and after applying CNG-SFDA. (b), (d): Examples of clean and noisy data
Effect on clean and noisy region. Fig. 3 shows the effectiveness of distinguishing clean and noisy data within clusters. From the gray boxes in (a) and (c) of Fig. 3, we observed that applying CNG-SFDA resulted in a reduction of noisy samples and clustering around clean samples. Additionally, as shown in (a), CNG-SFDA correctly identified noisy samples early in training, effectively utilizing them alongside clean samples. Interestingly, as seen in (d), samples correctly predicted to have noisy pseudo-labels were indeed difficult samples to classify.
Qualitative evaluation on distinguishing clean and noisy regions. Fig. 4 showcases samples with clean and noisy pseudo-labels predicted by CNG-SFDA at each target domain, along with the PL (Pseudo-Label) and CP (Clean Probability) of each sample. We observed that noisy images tend to have lower CP values, indicating difficulty in classification. These noisy images often contain objects outside the class or are challenging to perceptually classify. On the other hand, clean images mostly comprise easy samples containing only objects corresponding to the ground truth classes. This qualitative analysis confirms the reliability of CNG-SFDA in identifying clean and noisy images.
@@
Capability of Continual Test-Time Adaptation. Since CNG-SFDA works well in online SFDA (Test-time adaptation) settings, we conducted experiments on DomainNet-126 in continual TTA settings [52]. Continual TTA is particularly useful for scenarios where the target data distribution is not known in advance or is constantly changing. Consequently, the source model is dynamically adapted to a sequence of test domains in an online manner. For DomainNet-126, we randomly generated four different domain sequences, ensuring that each domain serves as the source domain once. Table. 7 reveals that CNG-SFDA surpasses SOTA [13] in the continual TTA setting, which demonstrates that CNG-SFDA robustly operates even in dynamic scenarios where unknown target domains continuously change.
@@
Robustness of hyper-parameters. As shown in Fig. 5, we investigate the robustness of CNG-SFDA versus the clean probability threshold $\alpha$ and $K$ in Eq. 14. In addition, we analyze the sensitivity of the memory queue length, which is the criterion for turning on and off soft voting for pseudo-labeling for online SFDA. These results demonstrate the robustness of CNG-SFDA to the hyper-parameters.
Linear vs Exponential weight. Table. 9 presents the effectiveness of the exponential weighting in Eq. 11, comparing linear one and without one. In particular, CNG-SFDA shows performance improvement relative to previous SFDA methods, even when linear weighting is utilized.
@@ 5 Conclusion
We introduced CNG-SFDA, a novel online-offline source-free domain adaptation approach in image classification. To ensure the comprehensive transfer of target domain knowledge to the source model, effective learning of clean and noisy pseudo-labels from the target domain is crucial. To address this, we propose to distinguish clean and noisy regions at the cluster level. Then, we apply different training strategies to each region by considering their respective characteristics. Our experiments and analysis prove the effectiveness of CNG-SFDA compared to other approaches.

### v3

@@ Abstract
Domain shift occurs when training (source) and test (target) data diverge in their distribution. Source-Free Domain Adaptation (SFDA) addresses this domain shift problem, aiming to adopt a trained model on the source domain to the target domain in a scenario where only a well-trained source model and unlabeled target data are available. In this scenario, handling false labels in the target domain is crucial because they negatively impact the model performance. To deal with this problem, we propose to update cluster prototypes (i.e., centroid of each sample cluster) and their structure in the target domain formulated by the source model in online manners. In the feature space, samples in different regions have different pseudo-label distribution characteristics affected by the cluster prototypes, and we adopt distinct training strategies for these samples by defining clean and noisy regions: we selectively train the target with clean pseudo-labels in the clean region, whereas we introduce mix-up inputs representing intermediate features between clean and noisy regions to increase the compactness of the cluster. We conducted extensive experiments on multiple datasets in online/offline SFDA settings, whose results demonstrate that our method, CNG-SFDA, achieves state-of-the-art for most cases. Code is available at https://github.com/hyeonwoocho7/CNG-SFDA.
@@
SFDA Clean-and-Noisy Region Online Prototype
@@ 1 Introduction
In this work, we introduce Clean-and-Noisy region Guided SFDA, namely CNG-SFDA, a novel SFDA approach that focuses on addressing pseudo-labels of samples near cluster boundaries given the domain knowledge from a source model. We first partition Clean and Noisy regions within clusters formed in the target domain by leveraging online-updating cluster prototype (i.e., the centroid of each cluster) to measure the distance between the prototypes and the unlabeled target samples. Then, distinct training strategies are adopted for the clean and noisy regions due to their divergent local characteristics.
For the clean region, we conduct selective training using reliable pseudo-labels derived from predictions of the nearest neighborhoods. In the noisy region, we leverage mix-up inputs that represent intermediate features in gray zone bridging clean and noisy regions in Fig. 1, unlike other methods that directly employ or discard the unreliable pseudo-labels for training. To draw these mix-up inputs toward the clean region effectively, we assign a “clean probability” to each sample indicating the likelihood of having clean pseudo-labels and introduce a mixed-clean probability defined from the clean probability of all target samples with clean and noisy, thereby enabling effective training on samples in noisy regions.
@@
  • We validate CNG-SFDA on various benchmarks for online/offline SFDA in single/multi-source settings, which empirically demonstrate the superiority of our approach to SOTA by a large margin in the challenging scenarios.
@@ 2.2 Online-Offline Source-Free Domain Adaptation
Recently, several methods [31, 51, 4, 25, 64, 33] have been proposed for handling the domain shift in online-offline source-free domain adaptation (SFDA) settings. SHOT [31] introduces a class centroid used for pseudo-labeling for offline SFDA. Tent [51] optimizes entropy minimization for online SFDA. CoWA-JMDS [28] utilizes the joint model-data structure as sample-wise weights to represent target domain knowledge effectively. CRCo [64] proposes a probability-based similarity between target samples by embedding the source domain class relationship.
@@
Figure 2: Framework of CNG-SFDA (a): partition the cluster into ‘Clean’ (close to the cluster prototype) and ‘Noisy’ (far from the cluster prototype) region based on the clean probability distribution. (b): training samples in clean regions with pseudo-labels generated from prediction of closest features. (c): reducing a gap between clean and noisy region with mix-up weight. (d): prototype/instance-aware contrastive learning.
@@ 3 The Proposed Method
In this work, we address Source-Free Domain Adaptation (SFDA) for the image classification task. In SFDA, we are given a source model trained on labeled source data $\mathcal{D}_{s}=\{x_{s}^{i},y_{s}^{i}\}_{i=1}^{N_{s}}$ and unlabeled target data $\mathcal{D}_{t}=\{x_{t}^{i}\}_{i=1}^{N_{t}}$, where $x_{s}^{i}$, $x_{t}^{i}$ represent the input images and $N_{s}$, $N_{t}$ denote the numbers of data in the source and target domains, respectively. We focus on the closed-set SFDA where the target domain shares the same $C$ classes as the source domain.
@@ 3.1 Distinguishing Clean and Noisy Regions in Clusters
$\small{\mu}_{k}=\text{norm}(\frac{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)g(x_{t}^{i})}{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)}),$ (4)
$\small{\sigma}_{k}=\frac{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)(g(x_{t}^{i})-\mu_{k})^{T}(g(x_{t}^{i})-\mu_{k})}{\sum_{i}p(\hat{y}_{t}^{i}=k|x_{t}^{i},\theta)}$ (5)
where $\text{norm}(\cdot)$ is l2-normalization such that $\left\|{\mu}_{k}\right\|_{2}$$=$$1$. With ${\mu}_{k},{\sigma}_{k}$ estimated from $f(\cdot)$, the Eq. (3) is defined by the probability density function:
@@ 3.2 Training Strategy for Clean and Noisy Regions
In preliminary study, we observed that clean and noisy regions within a cluster have divergent local characteristics. In the clean region, the samples, as well as their closest features, have reliable pseudo-labels $\hat{y}_{t}$ generated by Eq. 1, as they are also in the clean regions. On the other hand, the noisy regions contain samples with incorrect $\hat{y}_{t}$, as they are located near the cluster boundaries. To deal with this, we employ distinct training strategies for $\{x_{t}^{cr}\}_{cr=1}^{N_{cr}}$ and $\{x_{t}^{nr}\}_{nr=1}^{N_{nr}}$.
@@
Plus, we add a regularization promoting diverse predictions similar to [4]:
@@
where $\bar{p}_{t}=\mathbb{E}_{x_{t}\in\mathcal{X}_{t}}\sigma(f(T_{s}(x_{t})))$ and $L_{div}$ prevents the model from trusting false pseudo-labels.
@@ 3.3 Instance and Prototype-aware Contrastive Learning
Moreover, to learn instance-discriminate features, we leverage queries $q_{t}$ and keys $k_{t}=\bar{g}(T_{\bar{s}}(x_{t}))$ as the positive pair as in [18]. To build the negative pairs $\mathcal{N}_{p}$ of $q_{t}$, we store $k_{t}$ to the memory queue $Q_{k}=\{k_{t}^{i}\}_{i=1}^{M}$ with the length $M$, but $Q_{k}$ may have features which have the same class as the corresponding class $\hat{y}_{t}$ of $q_{t}$. In our work, to eliminate negative pairs with the same class, we utilize the nearest feature information in $Q_{k}$. We exclude the negative pairs when $\{\hat{y}_{t}^{n}\}_{n=1}^{N}$ of their $N$ closest features are at least one the same as $\hat{y}_{t}$, as shown in Fig. 2 (d). Our instance-aware contrastive loss reflecting the behavior above is given by:
@@ 4.1 Experimental Setup
4.2 Experimental Results
Table. 2, 4, and 4 show the performance of CNG-SFDA and the compared methods across 3 different datasets in the single-source setting. Table. 6 shows results in the multi-source setting. In each table, $\dagger$ indicates the results that we reproduced from the source code, and ‘SF’ and ‘On’ denote ‘source-free’ and ‘online’, respectively. All reported results are average accuracies from three seeds.
@@ 4.2 Experimental Results
Multi-Source DA. Table. 6 shows the performance of Multi-Source Unsupervised Domain Adaptation (MSUDA) on PACS. In MSUDA, each domain is treated as the target, while the source model is trained by aggregating all other domains without domain labels. In source-free MSUDA, CNG-SFDA achieves the highest accuracy by training pseudo-labels for the target with different strategies depending on whether the data is in clean or noisy regions, whereas SHOT++ [32] produces pseudo-labels using the centroid of the nearest cluster.
Method SF On A C P S Avg SIB [20] 88.9 89.0 98.3 82.2 89.6 T-SVDNET [30] 90.4 90.6 98.5 85.4 91.2 iMSDA[27] 93.7 92.4 98.4 89.2 93.4 Source Only - - 77.9 72.8 95.7 63.5 77.5 $\text{SHOT}^{\dagger}$ [31] 90.7 88.1 98.5 75.4 88.2 $+$ [32] 92.3 89.7 98.8 75.5 89.1 CNG-TTA 93.7 93.3 98.7 87.8 93.4 Table 5: Classification Accuracy (%) on PACS for Multi-Source Domain Adaptation (ResNet-18 Backbone). PL $L_{cr}$ $L_{ccp}$ $L_{div}$ $L_{inst}$ $L_{prt}$ VisDA-C DomainNet PACS 84.9 58.5 75.7 88.0 65.8 78.8 89.6 67.4 79.7 89.0 67.7 80.6 90.1 69.7 81.1 90.5 70.4 82.4 Table 6: The effectiveness of each component across all datasets is validated by classification accuracy (%).
[…]
Figure 3: Feature distribution of clean and noisy samples in target domain. Here, ‘clean’ and ‘noisy’ refer to cases where the Ground-Truth (GT) and the Pseudo-Label (PL) match and mismatch. Circle and triangular denote clean and noisy samples predicted by CNG-SFDA. For clean samples (circle), each color represent the GT class. For noisy samples (triangular), pink and black triangular represent cases where CNG-SFDA correctly (i.e., Prediction: Noisy, GT: Noisy) and incorrectly (i.e., Prediction: Noisy, GT: Clean) identifies noisy data.
@@ 4.3 Analysis and Discussion
Effect on clean and noisy region. Fig. 3 shows the effectiveness of distinguishing clean and noisy data within clusters. From the gray boxes in (a) and (c) of Fig. 3, we observed that applying CNG-SFDA resulted in a reduction of noisy samples and clustering around clean samples. Additionally, as shown in (a), CNG-SFDA correctly identified noisy samples early in training, effectively utilizing them with clean samples. Interestingly, as seen in (d), samples correctly predicted to have noisy pseudo-labels were indeed difficult samples to classify.
Capability of Continual Test-Time Adaptation. We also conducted experiments on DomainNet-126 in continual Test-time adaptation settings [52]. Continual TTA is particularly useful for scenarios where the target data distribution is not known in advance or is constantly changing. Consequently, the source model is dynamically adapted to a sequence of test domains in an online manner. For DomainNet-126, we randomly generated four different domain sequences, ensuring that each domain serves as the source domain once. Table. 7 reveals that CNG-SFDA surpasses SOTA [13] for the continual TTA, proving that its robustness in dynamic scenarios where target distributions are changing.
Qualitative evaluation on distinguishing clean and noisy regions. Fig. 4 showcases samples with clean and noisy pseudo-labels predicted by CNG-SFDA, along with the PL and CP of each sample. We observed that noisy images tend to have lower CP values, indicating difficulty in classification. These noisy images often contain objects outside the class or are challenging to perceptually classify. On the other hand, clean images mostly comprise easy samples containing only objects corresponding to the ground truth classes. This qualitative analysis confirms the reliability of CNG-SFDA in identifying clean and noisy images.
Robustness of hyper-parameters. Fig. 5 shows the robustness of CNG-SFDA versus the clean probability threshold $\alpha$ and $K$ in Eq. 14. In addition, we analyze the sensitivity of the memory queue length, which is the criterion for turning on-off soft voting for pseudo-labeling for online SFDA. These results demonstrate the robustness of CNG-SFDA to the hyper-parameters.
Linear vs Exponential weight. Table. 9 presents the effectiveness of the exponential weighting in Eq. 11, comparing linear one and without one. CNG-SFDA outperforms previous SFDA methods, even with linear weighting.
@@ 5 Conclusion
We introduced CNG-SFDA, a novel online-offline SFDA approach in image classification. To ensure the comprehensive transfer of target domain knowledge to the source model, effective learning of clean and noisy pseudo-labels from the target domain is crucial. To address this, we propose to distinguish clean and noisy regions at the cluster level. Then, we apply different training strategies to each region by considering their respective characteristics. Our experiments and analysis prove the effectiveness of CNG-SFDA compared to other approaches.
Acknowledgements This research was conducted with resources and endless support from VUNO Inc, and Won Hwa Kim was supported by Graduate School of AI at POSTECH (IITP-2019-0-01906).